38) Statistics. A survey of 400 people found that 66 were left-handed. What percent of those surveyed were left-handed?
Answer by tutorcecilia(2144)

Statistics:
Use the formula for percent:
Part/Whole=%/100
.
66/400= x%/100 [plug-in the values and solve for x]
(66)(100)=(400)(x%) [cross-multiply]
66000=400x% [solve for x]
66000/400=400x%/400
16.5=x%
16.5%=x%
.
Checking:
Part/Whole=%/100
66/100=16.5%/100
.165=.165 [checks out]